OpenStudy (yaelinapeartree):

What is the solution to the equation 3.6m − 2.7 = −1.8m? m = 0.25 m = 0.5 m = 1.25 m = 1.5

OpenStudy (erak):

What do you think you should do first?

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)

OpenStudy (yaelinapeartree):

Well, subtract 2.7 from 3.6?

OpenStudy (erak):

So "m" is the variable you are solving for, right? 3.6m and 2.7 cannot be subtracted because 2.7 is a constant whereas 3.6m has the variable attached to it. You would instead move all the numbers with "m" attached to one side of the equation and have the constant, 2.7, on the other.

OpenStudy (yaelinapeartree):

so, 3.6m - 1.8m - 2.7 = -1.8m ?

OpenStudy (yaelinapeartree):

well, 1.8m - 2.7 on the other side?

OpenStudy (erak):

So when you move a number over, it switches to its opposite sign. For example: 3.6m - 2.7 = -1.8m 3.6 m + 1.8m - 2.7 = 0 Does that kinda make sense?

OpenStudy (yaelinapeartree):

yes

OpenStudy (erak):

So now try moving the constant over to the right side, so all the terms on the left side have "m" attached

OpenStudy (yaelinapeartree):

3.6m + 1.8m = 2.7

OpenStudy (yaelinapeartree):

Right? I'm sorry, my brain is fried already but I've gotta finish this stuff

OpenStudy (erak):

yeah, and since 3.6m and 1.8m both have "m", you can add them together. Just add their coefficients (numbers in front of the m) and keep the m there.

OpenStudy (yaelinapeartree):

5.4m = 2.7

OpenStudy (yaelinapeartree):

so m = .5 ?

OpenStudy (erak):

yeah, congrats

OpenStudy (yaelinapeartree):

Thank you:)